You are designing a circular unbanked freeway ramp with a radius of 100. m, the coefficient of static friction between the tires and asphalt is 0.800 dry and 0.400 wet. Find the maximum speed that a car can travel through the ramp during when the road is dry and wet in mph (miles per hour).

Vmax, dry =

Vmax, wet =
Answer #1

speed during the turn = sqrt(urg)

where u = friction

r = radius

g= acceleration due to gravity

v(max , dry) = sqrt(0.8*100*9.8)

= 28 m/s

v(max , wet) = sqrt(0.4*100*9.8) = 19.8 m/s